OpenStudy (anonymous):

example of ordinary differential equation and partial differential equation.

OpenStudy (anonymous):

Hi nice and Your pic shows You are Passionate about photography........ nice pic ;) Here's and URL for the question above presented by KHAN ACADEMY : http://www.khanacademy.org/math/differential-equations/v/what-is-a-differential-equation

OpenStudy (anonymous):

Well he explained it. A differential equation is non-linear if the function or its derivative (of any order) is multiplied to itself or to any of its derivatives. e.g. x.(d^2y/dx^2)+dy/dx+5=0 is linear x.(d^2y/dx^2)^2=sin(x) is non-linear since the 2nd derivative is multiplied to itself. y.(d^2y/dx^2)=sin(x) is non-linear since the 2nd derivative is multiplied to the function y.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

id like to ask which is the partial and the ordinary?

OpenStudy (inkyvoyd):

ODE: y'+y''-y+x=0

OpenStudy (inkyvoyd):

PDE: involves the  or f(blah,blah)

OpenStudy (inkyvoyd):

a layman's view: ODE are hard. PDE are harder.

OpenStudy (unklerhaukus):

inkyvoyd i think you got that backwards \[y_x=\partial_xy=\frac{\partial y}{\partial x}\]